# Robust mass lumping and outlier removal strategies in isogeometric analysis

### v1

@@ 5.2 Multipatch geometries
In this example, the Laplace eigenvalue problem with pure Neumann boundary conditions is solved on a shifted and rotated square, as shown in Figure 5.15. The physical domain is embedded in a regular unit square, which is trimmed to fit the domain’s boundaries. We test our block lumping strategies, combined with a rank $80$ scaling, and compare them to the standard row-sum technique. System matrices for trimmed domains generally do not feature a Kronecker structure, not even in their sparsity. However, by padding them with zero entries, they may be embedded in larger matrices whose structure allows applying the block lumping operator. Once the operator has been applied, they are trimmed back to their original size by removing the artificial rows and columns. Thus, our lumping strategies also apply to trimmed geometries. The spectrum (after discarding the zero eigenvalue) is reported in Figures 15(a) and 15(b) for $p=2$ and $p=3$, respectively. It is well-known that small trimmed elements heavily deteriorate the conditioning of the system matrices (among other issues) [42, 62]. Beyond a certain threshold, the computations altogether are no longer accurate. Our rotation angle, while still being unfavorable, avoids such extreme situations. Nevertheless, one should exercise caution when computing eigenvalues of matrix pairs $(A,B)$ with a heavily ill-conditioned matrix $B$. In our numerical experiments, we have computed the eigenvalues of the equivalent matrix pair $(DAD,DBD)$, where $D=\operatorname{\operatorname{diag}}(d_{1},\dots,d_{n})$ with $d_{i}=1/\sqrt{b_{ii}}$ for $i=1,\dots,n$ is a Jacobi preconditioner for $B$ [62]. In the context of trimming, the conditioning of $DBD$ is generally orders of magnitude better than the one of $B$, which improves the stability of eigensolvers.
@@
As shown in Figures 15(a) and 15(b), the outlier eigenvalues for this problem are particularly pronounced, as evidenced by the sharp change of curvature in the spectrum. Although the row-sum technique is most effective at improving the CFL condition, it is also most inaccurate, even for moderate spline degrees. Our block lumping method drastically improves the accuracy but also leads to more restrictive CFL conditions. However, combining it with scaling strongly mitigates this effect. The fast increase of the outlier eigenvalues also speeds up the convergence of the Lanczos method, which in turn improves the efficiency of the method for explicit dynamics.
@@ 6 Conclusion
In this article, we have proposed robust mass lumping and outlier removal techniques for nontrivial isogeometric discretizations, including multipatch and trimmed geometries. Our mass lumping techniques provably do not deteriorate the CFL condition of the original problem and oftentimes improve it, thereby extending the methods proposed in [32] to more realistic settings. For a significant increase in step size, we suggest purging persistent outliers by deflating the spectrum. Contrary to existing outlier removal techniques, this method only relies on standard eigensolvers, whose cost per iteration is comparable to explicit time integration methods and whose efficiency is enhanced by the large eigenvalue gaps characterizing outliers. Numerical experiments have shown that the cost for computing a few eigenpairs is rapidly amortized by the subsequent increase in critical time step.
Acknowledgments
The second author kindly acknowledges the support of the SNSF through the project “Smoothness in Higher Order Numerical Methods” n. TMPFP2_209868 (SNSF Swiss Postdoctoral Fellowship 2021).
The third author kindly acknowledges the support of SNSF through the project “PDE tools for analysis-aware geometry processing in simulation science” n. 200021_215099.
@@ Appendix A Appendix
It was shown in [4] that maximally smooth spline spaces provide better approximation per degree of freedom than $C^{0}$ finite element spaces in almost all cases of practical interest. In this section we shall see how this improved approximation guarantees that there are fewer highly inaccurate outlier modes in the case of smooth spline approximations than for $C^{0}$ FEM for any elliptic PDE. In this section we use the standard Sobolev spaces
@@
see [63, Chapter 6.3] and [64, Chapter 3.2] for the details. In fact, for any nonnegative integer $m$, if the coefficients in $a(\cdot,\cdot)$ are $C^{m+1}$ and the boundary $\partial\Omega$ is $C^{m+2}$ then we have the estimate [63, Chapter 6.3]
@@
Following [5] we define the constant $C_{p,k,r}$ as follows. If $k=p-1$, we let
@@
where the constant $C_{\text{Geo}}$ only depends on the geometry maps $F_{i}$, $i=1,\ldots,N_{p}$ and is explicitly given in [5]. In classical finite element methods the smoothness $k=0$ and in this case the constant $C_{p,0,r}$ satisfies, for $p\geq 2$, the following inequalities [5, Remark 3]
@@
It follows from the spline dimension formula that the mesh size $h\sim d(p-k)/n$ and thus $C_{p,k,r}h^{r}\sim C_{p,k,r}(d(p-k)/n)^{r}$. More explicitly, with the upper bounds previously obtained:
@@
    $C_{p,k,r}\frac{d(p-k)}{n}\lesssim\begin{cases}\left(\frac{d}{2n}\right)^{r}&\text{if }k=0,\\ \left(\frac{d}{\pi n}\right)^{r}&\text{if }k=p-1.\end{cases}$
@@
    $C_{p,k,r}\frac{d(p-k)}{n}\lesssim\begin{cases}\left(\frac{ed}{4n}\right)^{r}&\text{if }k=0,\\ \left(\frac{d}{\pi n}\right)^{r}&\text{if }k=p-1.\end{cases}$
Although these estimates are only upper bounds, the difference between the $C^{0}$ and $C^{p-1}$ cases is readily appreciated. For instance, for $p=3$ and $r=4$ the upper bound is about $20$ times smaller in the $C^{p-1}$ case than in the $C^{0}$ case. The reader may refer to [5, Figs. 1 and 2] for a graphical comparison of the constants for different values of $p$ and $r$. In fact, by using a lower bound on the best approximation constant in the $C^{0}$ case it is shown in [4] that the maximally smooth approximation constant $C_{p,p-1,r}$ becomes exponentially better than the best achievable approximation constant for $k=0$ as the degree $p\geq 3$ and $r$ increase.
@@
From our previous discussion, it follows that for fixed $n$ and $r$ and for a given tolerance, maximally smooth splines allow to approximate a larger fraction of the eigenfunctions (and eigenvalues) than $C^{0}$ finite element spaces. Moreover, if the PDE satisfies elliptic regularity then the error estimate in (A.9) can be further improved.
@@
The improved approximation for maximally smooth splines compared with $C^{0}$ finite element spaces will only get better as $r$ increases in Theorem A.2, and we are guaranteed good approximation of a larger fraction of the eigenfunctions and eigenvalues than for $C^{0}$ FEM.

### v2

@@ 5.3 Trimmed geometries
In this example, the Laplace eigenvalue problem with pure Neumann boundary conditions is solved on a shifted and rotated square. The physical domain is embedded in a regular unit square discretized with $N=80$ subdivisions in each direction and trimmed to fit the domain’s boundaries. Figure 5.15 provides an illustration for a coarser mesh with only $20$ subdivisions in each direction. We test our block lumping strategies, combined with a rank $80$ scaling, and compare them to the standard row-sum technique for quadratic and cubic order spline discretizations. System matrices for trimmed domains generally do not feature a Kronecker structure, not even in their sparsity. However, by padding them with zero entries, they may be embedded in larger matrices whose structure allows applying the block lumping operator. Once the operator has been applied, they are trimmed back to their original size by removing the artificial rows and columns. Thus, our lumping strategies also apply to trimmed geometries. The resulting system size was $1244$ and $1321$ for $p=2$ and $p=3$, respectively. The scaled and unscaled spectra of $(\mathcal{K},\mathcal{M})$, $(\mathcal{K},P_{1})$ and $(\mathcal{K},\mathcal{P}_{i})$ for $i=1,2$ is reported in Figures 16(a) and 16(b) for $p=2$ and $p=3$, respectively, after discarding the zero eigenvalue. It is well-known that small trimmed elements heavily deteriorate the conditioning of the system matrices (among other issues) [43, 66]. Beyond a certain threshold, the computations altogether are no longer accurate. Our rotation angle, while still being unfavorable, avoids such extreme situations. Nevertheless, one should exercise caution when computing eigenvalues of matrix pairs $(A,B)$ with a heavily ill-conditioned matrix $B$. In our numerical experiments, we have computed the eigenvalues of the equivalent matrix pair $(DAD,DBD)$, where $D=\operatorname{\operatorname{diag}}(d_{1},\dots,d_{n})$ with $d_{i}=1/\sqrt{b_{ii}}$ for $i=1,\dots,n$ is a Jacobi preconditioner for $B$ [66]. In the context of trimming, the conditioning of $DBD$ is generally orders of magnitude better than the one of $B$, which improves the stability of eigensolvers.
@@
As shown in Figures 16(a) and 16(b), the outlier eigenvalues for this problem are particularly pronounced, as evidenced by the sharp change of curvature in the spectrum. Although the row-sum technique is most effective at improving the CFL condition, it is also most inaccurate, even for moderate spline degrees, as highlighted in [40]. Our block lumping method drastically improves the accuracy but also leads to more restrictive CFL conditions. However, combining it with scaling strongly mitigates this effect. The fast increase of the outlier eigenvalues also speeds up the convergence of the Lanczos method, which in turn improves the efficiency of the method for explicit dynamics. Indeed, for $p=2$ and a tolerance of $10^{-3}$, Lanczos needed $486$ iterations for the row-sum technique $P_{1}$ while it needed only $243$ iterations for the block lumped mass matrices $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. The number of iterations rose to $891$ for the row-sum while it remained constant for $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ when increasing the spline degree to $p=3$. The widely different behavior of the outliers for the row-sum technique and their block counterpart explains the stunning difference in iteration count. While the outliers for $\mathcal{P}_{i}$ sharply increase as they approach those for the consistent mass, those for $P_{1}$ barely stand out and Lanczos struggles to converge to these “outliers”. The spectra for $40$ uniformly spaced rotation angles between 00 and $2\pi$ are shown in Figures 17(a) and 17(b) for $p=2$ and $p=3$, respectively, and consistently capture the same trend.
@@
(a) Spectrum for $p=2$
(b) Spectrum for $p=3$
Figure 5.17: Spectrum for $40$ uniformly spaced rotation angles between 00 and $2\pi$
@@ 6 Conclusion
In this article, we have proposed mass lumping and outlier removal techniques for nontrivial isogeometric discretizations, including multipatch and trimmed geometries. Our mass lumping techniques focus on reducing the bandwidth of the consistent mass and therefore significantly accelerate sparse direct solvers. Moreover, they provably do not deteriorate the CFL condition of the original problem and oftentimes improve it, thereby extending the methods proposed in [32] to more realistic settings. For a significant increase in step size, we suggest purging persistent outliers by deflating the spectrum. This outlier removal technique is independent of the lumping strategies proposed herein and contrary to other existing techniques, it only relies on standard eigensolvers, whose cost per iteration is comparable to explicit time integration methods and whose efficiency is enhanced by the large eigenvalue gaps characterizing outliers. Numerical experiments have shown that the cost for computing a few eigenpairs is rapidly amortized by the subsequent increase in critical time step.
@@ Appendix A Appendix
It was shown in [4] that maximally smooth spline spaces provide better approximation per degree of freedom than $C^{0}$ finite element spaces in almost all cases of practical interest. In this section we shall see how this improved approximation guarantees that there are fewer highly inaccurate outlier eigenvalues in the case of smooth spline approximations than for $C^{0}$ FEM for any elliptic PDE. In this section we use the standard Sobolev spaces
@@
see [67, Chapter 6.3] and [68, Chapter 3.2] for the details. In fact, for any nonnegative integer $m$, if the coefficients in $a(\cdot,\cdot)$ are $C^{m+1}$ and the boundary $\partial\Omega$ is $C^{m+2}$ then we have the estimate [67, Chapter 6.3]
@@
Following [5], we define the constant $C_{p,k,r}$ as follows. If $k=p-1$, we let
@@
where the constant $C_{\text{Geo}}$ only depends on the geometry maps $F_{i}$, $i=1,\ldots,N_{p}$ and is explicitly given in [5]. In classical finite element methods, the smoothness $k=0$ and in this case the constant $C_{p,0,r}$ satisfies, for $p\geq 2$, the following inequalities [5, Remark 3]
@@
It follows from the spline dimension formula that the mesh size $h\sim(p-k)/n^{1/d}$ and thus $C_{p,k,r}h^{r}\sim C_{p,k,r}((p-k)/n^{1/d})^{r}$. More explicitly, with the upper bounds previously obtained:
@@
    $C_{p,k,r}\left(\frac{p-k}{n^{1/d}}\right)^{r}\lesssim\frac{1}{n^{r/d}}\begin{cases}\left(\frac{1}{2}\right)^{r}&\text{if }k=0,\\ \left(\frac{1}{\pi}\right)^{r}&\text{if }k=p-1.\end{cases}$
@@
    $C_{p,k,r}\left(\frac{p-k}{n^{1/d}}\right)^{r}\lesssim\frac{1}{n^{r/d}}\begin{cases}\left(\frac{e}{4}\right)^{r}&\text{if }k=0,\\ \left(\frac{1}{\pi}\right)^{r}&\text{if }k=p-1.\end{cases}$
Although these estimates are only upper bounds, the difference between the $C^{0}$ and $C^{p-1}$ cases is readily appreciated. For instance, for $p=3$ and $r=4$ the upper bound is about $20$ times smaller in the $C^{p-1}$ case than in the $C^{0}$ case. The reader may refer to [5, Figs. 1 and 2] for a graphical comparison of the constants for different values of $p$ and $r$. In fact, by using a lower bound on the best approximation constant in the $C^{0}$ case, it is shown in [4] that the maximally smooth approximation constant $C_{p,p-1,r}$ becomes exponentially better than the best achievable approximation constant for $k=0$ as the degree $p\geq 3$ and $r$ increase.
@@
From our previous discussion, it follows that for fixed $n$ and $r$ and for a given tolerance, maximally smooth splines allow to approximate a larger fraction of the eigenvalues than $C^{0}$ finite element spaces. Moreover, if the PDE satisfies elliptic regularity then the error estimate in (A.9) can be further improved.
@@
The improved approximation for maximally smooth splines compared with $C^{0}$ finite element spaces will only get better as $r$ increases in Theorem A.2, and we are guaranteed good approximation of a larger fraction of the eigenvalues than for $C^{0}$ FEM. Using the arguments presented in [69], the statement is also valid for eigenfunctions.